### v2

# Think Twice: Branch-and-Rethink Reasoning Reward Model

Abstract
Large language models (LLMs) increasingly rely on thinking models that externalize intermediate steps and allocate extra test-time compute, with think-twice strategies showing that a deliberate second pass can elicit stronger reasoning. In contrast, most reward models (RMs) still compress many quality dimensions into a single scalar in one shot, a design that induces judgment diffusion: attention spreads across evaluation criteria, yielding diluted focus and shallow analysis. We introduce branch-and-rethink (BR-RM), a two-turn RM that transfers the think-twice principle to reward modeling. Turn 1 performs adaptive branching, selecting a small set of instance-critical dimensions (such as factuality and safety) and sketching concise, evidence-seeking hypotheses. Turn 2 executes branch-conditioned rethinking, a targeted reread that tests those hypotheses and scrutinizes only what matters most. We train with GRPO-style reinforcement learning over structured two-turn traces using a simple binary outcome reward with strict format checks, making the approach compatible with standard RLHF pipelines. By converting all-at-once scoring into focused, second-look reasoning, BR-RM reduces judgment diffusion and improves sensitivity to subtle yet consequential errors while remaining practical and scalable. Experimental results demonstrate that our model achieves state-of-the-art performance on three challenging reward modeling benchmarks across diverse domains. The code and the models will be released soon.
Source Code: https://github.com/yzjiao/BR-RM.
Model Checkpoints: nvidia/Qwen3-Nemotron-14B-BRRM and nvidia/Qwen3-Nemotron-8B-BRRM
[…]
1 Introduction
Large language models (LLMs) increasingly solve hard problems by reasoning rather than merely responding (Shao et al., 2024; DeepSeek-AI et al., 2025). Under the banner of “reasoning LMs,” models externalize intermediate steps – e.g., chain-of-thought, scratchpads, and reflect-revise traces – and allocate more test-time compute to difficult instances (Wei et al., 2022; Nye et al., 2021; Shinn et al., 2023; Yao et al., 2023; Wang et al., 2023). These procedures convert raw capacity into usable reasoning: they surface hypotheses, check local inferences, and expose mistakes that would otherwise remain hidden. Recent “thinking” models and analyses of test-time compute further highlight the benefits of deliberate, multi-step inference and adaptive allocation (Snell et al., 2024).
Yet a single explicit-reasoning pass is often brittle – a quiet factual slip, a local contradiction, or a small logic bug can persist because attention is spread across many potential concerns. Intuitively and empirically, thinking twice helps: a second, targeted pass re-reads with hypotheses, directs scrutiny to suspicious spans, and reallocates compute where uncertainty is highest (Wang et al., 2023; Tian et al., 2025). This naturally raises the question: if solvers benefit from a second look, should our judges – that grade model outputs – be designed to do the same?
Enter Reward Models (RMs), the workhorse that turns preference signals into scalar rewards in RLHF- and RLAIF-style pipelines (Ouyang et al., 2022; Bai et al., 2022b). Scalar RMs typically output a single holistic score per response – implicitly aggregating all aspects of response quality in a single forward pass. Since the judge is asked to consider everything at once, it is discouraged from deep scrutiny of what matters most for a given instance.
To address this limitation, Generative Reward Models (GenRMs) ask the judge to explain before scoring: the RM first produces a short critique or rationale and then outputs a number, improving transparency and often robustness (Mahan et al., 2024; Zhang et al., 2025; Ye et al., 2025). Still, most GenRMs reason under a fixed rubric and ultimately collapse their critique into a single global decision, so the underlying all-at-once pressure remains. Based on this, Reasoning Reward Models (ReasonRMs) treat judgment itself as a reasoning task, deliberately allocating test-time compute prior to scoring (Chen et al., 2025b; Guo et al., 2025). Yet ReasonRMs typically reason broadly (rubric-wide) and do not enforce instance-adaptive focus or a second pass conditioned on discovered issues. Thus, even as they add reasoning, they do not guarantee that attention concentrates where risk is highest.
Figure 1: Comparing token allocation between our method and a recent ReasonRM (Guo et al., 2025) on two subsets from RM-Bench. Our method adaptively focuses its generative analysis on a few critical dimensions for each task (e.g., Information Accuracy for chat), while the baseline spreads its tokens broadly across many criteria. More details in Appendix 3.
We highlight a recurring weakness across existing RMs, which we call judgment diffusion. When forced to grade responses holistically, these models spread their attention thin across many possible criteria. The result is focus dilution, where no single dimension is examined with sufficient depth, and shallow analysis, where judgments skim the surface instead of probing concrete failure modes. Empirical studies reinforce this view: RMs and LLM judges often miss subtle factual slips or localized reasoning errors, and their decisions are easily swayed by stylistic factors rather than substantive quality (Lambert et al., 2025a; Liu et al., 2025b). To empirically ground our hypothesis, we conduct a preliminary study of the generation from a recent ReasonRM (Guo et al., 2025) on two challenging subsets of RM-Bench (Liu et al., 2025b). We analyze how this model allocates its generative token budget across different evaluation dimensions. Figure 1 shows this model spreads its analysis thinly across multiple dimensions without detailed reasoning on the most critical dimensions.
Motivated by these observations, we introduce a two-turn reward modeling framework that builds focus and rethinking into the RM. Instead of grading everything at once, our framework recasts judgment as branch-and-rethink: the model first narrows its focus to a few instance-critical dimensions, then performs a second, conditioned pass that scrutinizes only what matters most. Specifically, the first turn is Adaptive Branching. The RM selects a few relevant cognitive dimensions from nine candidates (e.g., factual accuracy, reasoning quality, safety) and produces a concise issue sketch pointing to potential weaknesses. This focuses attention where risk is highest, replacing diffusion with targeted scrutiny. The second turn is Branch-Conditioned Rethinking. Using the first-turn findings, the RM re-reads the response through the lens of flagged dimensions – e.g., verifying facts, checking brittle reasoning, or examining localized bugs. This issue-driven pass avoids the common failure mode of broad but shallow reasoning. For training, we adopt GRPO-style reinforcement learning to support two-turn traces: both turns produce structured outputs, and optimization uses a simple binary outcome reward with strict format checks. This keeps supervision clean, interoperates with standard RLHF infrastructure, and ensures compute is spent where a second look matters.
Extensive experiments on reward modeling benchmarks show that our method, BR-RM, achieves overall state-of-the-art performance across RewardBench (Lambert et al., 2025b), RMBench (Liu et al., 2025b), and RMB (Zhou et al., 2025), outperforming strong baselines across multiple domains, including reasoning, general knowledge, safety, and alignment with human preferences. Beyond final performance, we conduct extensive empirical analyses of BR-RM, including the ablation study of our training recipes, the detailed analysis of the reward designs, and concrete impacts of training data sources. These insights provide a deeper understanding of reward reasoning processes and will hopefully inspire the development of future reasoning reward models.
2 Related Work
2.1 Reasoning Language Models
Reasoning has emerged as a defining capability of modern LLMs, marking a shift from surface-level response generation to structured problem-solving. Early advances such as chain-of-thought prompting, scratchpads, tree-structured exploration, and self-consistency demonstrate that explicitly externalizing intermediate steps improves accuracy and interpretability in multi-step tasks (Wei et al., 2022; Nye et al., 2021; Yao et al., 2023; Wang et al., 2023). By surfacing hypotheses and local inferences, these methods transform raw capacity into usable reasoning, enabling models to identify and correct mistakes that would otherwise remain hidden. Beyond single-pass prompting, reflect–revise and multi-pass paradigms show that “thinking twice” helps models correct quiet factual slips and local logic bugs through targeted second looks (Shinn et al., 2023; Snell et al., 2024; Tian et al., 2025). Domain-focused reasoning systems (e.g., math) further highlight that deliberate computation and reinforcement learning can incentivize deeper reasoning traces (Shao et al., 2024; DeepSeek-AI et al., 2025). Collectively, these works suggest that allocating test-time compute and conditioning a second pass on suspected weaknesses can meaningfully improve reliability. We draw on this intuition, transferring the “think twice” principle from solvers to judges by enforcing a focused, issue-conditioned second pass in our framework.
2.2 Reward Models
Reward models (RMs) are central to alignment pipelines such as RLHF, converting human or AI preference signals into scalar rewards for optimization (Ouyang et al., 2022; Bai et al., 2022b). Classical discriminative RMs predict a single holistic score per response, implicitly aggregating in a single forward pass factuality, reasoning quality, safety, coding correctness, instruction following, style, etc. (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Bai et al., 2022a; Sun et al., 2023; Liu et al., 2024a; Zhu et al., 2024). While simple and scalable, this all-at-once design dilutes attention across dimensions, limiting sensitivity to subtle but consequential errors.
To improve transparency and robustness, Generative Reward Models (GenRMs) introduce explain-then-score: the model produces a short critique or rationale prior to emitting a score (Mahan et al., 2024; Zhang et al., 2025; Ankner et al., 2024; Liu et al., 2025a). However, most systems still collapse critiques to a single global decision and rarely enforce instance-adaptive focus. Recent Reasoning Reward Models (ReasonRMs) treat judgment itself as a reasoning task, allocating test-time compute prior to scoring (Chen et al., 2025b; Guo et al., 2025; Whitehouse et al., 2025). Yet many reason broadly over fixed rubrics, without guaranteeing that scrutiny concentrates on the instance-critical dimensions discovered during evaluation. DeepSeek-GRM (Liu et al., 2025c) prompts the model to generate instance-specific evaluation criteria, but does so in a single-turn, free-form manner and still maintains a fixed general rubric. In contrast, we found our structured two-turn “branch-and-rethink” approach to be crucial to reduce judgment diffusion and increase accuracy. Similar in spirit to our work, EvalPlanner (Saha et al., 2025) trains a judge model that first plans detailed evaluation steps before executing on them: the generated plan, however, is entirely open-ended – while we constrain the model to choose among specific criteria to focus on – and the approach relies on iterative Direct Preference Optimization (Rafailov et al., 2023) – while we use online Reinforcement Learning.
Parallel progress is the curation of preference data for training and evaluation. Large-scale training datasets, such as HelpSteer3 (Wang et al., 2025) and Skywork-Reward (Liu et al., 2024b, 2025a), provide high-quality preference signals that enable the training of more nuanced RMs. Other comprehensive benchmarks, such as RewardBench and RM-Bench (Lambert et al., 2025a; Liu et al., 2025b; Tan et al., 2025), systematically evaluate the ability of current models. Key findings from this line of work reveal low sensitivity to subtle content flips and a strong susceptibility to superficial heuristics, including length and position biases (Hu et al., 2024; Shi et al., 2024; Chen et al., 2024).
3 Method
Our approach, branch-and-rethink, reframes reward modeling from a direct regression task over a scalar value to a structured reasoning and generation task. We introduce a framework that compels the reward model to produce an explicit, structured deliberation trace. This is achieved through a sequential, two-stage generation process designed to mitigate the judgment diffusion observed in single-pass evaluators, by first establishing a focused analytical scope and then performing a deep, conditioned analysis. Fig. 2 contrasts our method with generative RMs and general Reasoning RMs.
3.1 Task Formulation
The standard reward modeling task is to learn a preference function from a dataset $\mathcal{D}=\{(x_{i},y_{i,1},y_{i,2},z_{i})\}_{i=1}^{N}$. For a given prompt $x_{i}$, the goal is to correctly identify the preferred response between two candidates, $y_{i,1}$ and $y_{i,2}$, as indicated by a human label $z_{i}\in\{1,2\}$.
[…]
3.2 The Branch-and-Rethink Framework
Figure 2: Illustration of our proposed method, comparing with GenRM and ReasonRM. Our Branch-and-Rethink Reasoning Reward Model first performs Turn 1: Adaptive Branching, where it selects a few critical dimensions (e.g., Logical Reasoning, Computation Precision) to focus its evaluation and detect specific issues. This focused analysis then informs Turn 2: Branch-Conditioned Rethinking, where the model conducts a deeper, issue-driven re-thinking to arrive at a final reward judgment, which is then used for reinforcement learning.
To produce the reasoning trace $\tau$, we employ a two-stage generative workflow. This architecture is the core of our method, designed to systematically address the failure modes of focus dilution and shallow analysis. The complete trace is a concatenation of the outputs from each stage: $\tau=\tau_{1}\circ\tau_{2}$.
3.2.1 Stage 1: Adaptive Branching for Focused Evaluation
The first stage directly targets focus dilution. To prevent the model’s evaluation from becoming diffuse, we constrain it to first select a small subset of critical evaluation criteria, $\mathcal{C}_{\text{sel}}$, from a predefined universal set of criteria, $\mathcal{C}$. For example, $\mathcal{C}$ could be {Factual Accuracy, Logical Coherence, Instruction Adherence, Creativity…}. The detailed evaluation criteria can be found in Appendix C.
This selection forces the model to hypothesize the most salient dimensions for the given comparison. Conditioned on this selection, the model also generates a preliminary analysis, $\alpha_{j}$, for each response $y_{j}$:
$(\mathcal{C}_{\text{sel}},\alpha_{1},\alpha_{2})\sim\pi_{\theta}(\cdot\mid x,y_{1},y_{2})$ (1)
The output of this stage, which includes the selected criteria and the initial analysis, constitutes the first part of the trace, $\tau_{1}=(\mathcal{C}_{\text{sel}},\alpha_{1},\alpha_{2})$.
3.2.2 Stage 2: Conditioned Rethinking for Deep Analysis
The second stage is designed to prevent shallow analysis. With the analytical scope established by $\tau_{1}$, the model performs a targeted re-evaluation. This Branch-Conditioned Rethinking is not an independent assessment but a deliberate critique of the responses through the lens of the criteria identified in Stage 1. To ensure this deliberation is principled, we incorporate task-specific evaluation hierarchies, $\mathcal{H}_{\text{task}}$, which provide a structured inductive bias. These hierarchies encode a stable ordering of importance for different criteria, such as:
  • Accuracy-Critical Tasks: $\text{Correctness}>\text{Methodology}>\text{Clarity}$
  • Creative Tasks: $\text{Intent Alignment}>\text{Quality}>\text{Novelty}$
The model is conditioned on $\tau_{1}$ and guided by $\mathcal{H}_{\text{task}}$ to generate the second part of the trace, $\tau_{2}$, which contains a detailed comparative judgment and the final decision, $\hat{z}$.
$\tau_{2}\sim\pi_{\theta}(\cdot\mid x,y_{1},y_{2},\tau_{1})$ (2)
The final verdict, $\hat{z}$, is then extracted from $\tau_{2}$.
3.3 Learning via Reinforcement Learning
3.3.1 The GRPO Training Algorithm
We employ Generalized Reward Policy Optimization (GRPO, Shao et al. 2024), a variant of PPO, chosen for its training stability and sample efficiency. The training loop (Algorithm 1) optimizes the policy $\pi_{\theta}$ to maximize a composite reward function over the distribution of generated traces.
The GRPO objective balances reward maximization with training stability via a clipped surrogate objective and a KL-divergence penalty against a reference policy $\pi_{\text{ref}}$:
$L_{\text{GRPO}}=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\min\left(\rho(\tau)A(\tau),\text{clip}(\rho(\tau),1-\epsilon,1+\epsilon)A(\tau)\right)\right]-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$ (3)
where $\rho(\tau)=\frac{\pi_{\theta}(\tau)}{\pi_{\text{ref}}(\tau)}$ is the probability ratio. Here $A(\tau)$ is a group-relative, critic-free advantage: for each prompt we sample $K$ traces, compute a terminal reward $r(\tau)$ per trace, subtract the per-prompt mean and divide by the per-prompt standard deviation, and use this centered (whitened) value as $A(\tau)$. We assign this scalar uniformly to all tokens in the trace.
Algorithm 1 Two-Turn RL Training with GRPO
1:for each training step do
2:  $B\leftarrow\text{SampleBatch}(\mathcal{D},\text{batch\_size})\triangleright$ Sample a batch of data
3:  $U\leftarrow\emptyset\triangleright$ Initialize an empty update buffer
4:  for each $(x_{i},y_{i,1},y_{i,2},z_{i})\in B$ do
5:   $\tau_{i,1}\leftarrow\pi_{\theta}(\cdot\mid x_{i},y_{i,1},y_{i,2})\triangleright$ Generate the branching trace
6:   $r_{1,i}\leftarrow\text{CalculateReward}(\tau_{i,1})\triangleright$ Calculate format reward for turn 1
7:   $U\leftarrow U\cup\{(x_{i},y_{i,1},y_{i,2},\tau_{i,1},r_{1,i})\}\triangleright$ Add first turn to buffer with reward $r_{1,i}$
8:   $\tau_{i,2}\leftarrow\pi_{\theta}(\cdot\mid x_{i},y_{i,1},y_{i,2},\tau_{i,1})\triangleright$ Generate the rethinking trace
9:   $r_{2,i}\leftarrow\text{CalculateReward}(z_{i},\tau_{i,2})\triangleright$ Calculate reward based on the final decision
10:   $U\leftarrow U\cup\{(x_{i},y_{i,1},y_{i,2},\tau_{i,1},\tau_{i,2},r_{2,i})\}\triangleright$ Add second turn to buffer with reward $r_{2,i}$
11:  end for
12:  $\text{UpdatePolicy}(\pi_{\theta},U)\triangleright$ Update policy using the GRPO objective after sampling K traces
13:end for
3.3.2 Reward Function Design
Our reward $R(\tau)$ aims to guide the learning process through a curriculum, by summing two terms:
$R(\tau)=R_{\text{format}}(\tau)+wR_{\text{outcome}}(\tau),$ (4)
where $w$ is the weight to balance two terms.
Format Reward: To enforce the two-stage format, we heavily penalize any format deviation.
$R_{\text{format}}(\tau)=\begin{cases}0&\text{if }\tau\text{ is well-formed}\\ \lambda_{\text{format}}&\text{if }\tau\text{ is malformed}\end{cases}$ (5)
The penalty constant $\lambda_{\text{format}}<0$ ensures that learning the correct generative format is prioritized.
Binary Outcome Reward: The substantive quality of a trace is rewarded based on the correctness of its final decision. This reward is only applied if the trace is structurally valid, i.e. $R_{\text{format}}(\tau)=0$.
$R_{\text{outcome}}(\tau)=\begin{cases}0&\text{if }\hat{z}=z\\ -1&\text{if }\hat{z}\neq z\end{cases}$ (6)
This binary reward (calculated only after the full trace is generated) encourages the model to produce a reasoning that leads to the correct outcome. The same reward $R(\tau)$ is assigned separately to the first turn (generation of the branching trace $\tau_{1}$) and the second turn (generation of the rethinking trace $\tau_{2}$ conditioned on $\tau_{1}$). This formulation makes it easy to integrate into standard implementations of the GRPO objective without requiring explicit support for multi-turn Reinforcement Learning.
4 Experiment
4.1 Experimental Setup
Training and Evaluation Data. To train the proposed model, we used a diverse mix of preference datasets, including HelpSteer3 (Wang et al., 2025), Skywork Reward Preference-80K (Liu et al., 2024b), Code-Preference-Pairs (Vezora, 2024), and Math-Step-DPO-10K (Lai et al., 2024). Following Chen et al. (2025b), model performance was measured by preference accuracy on RewardBench (Lambert et al., 2025b), RM-Bench (Liu et al., 2025b) and RMB (Zhou et al., 2025).
Baselines. Still following Chen et al. (2025b), the baselines are chosen to cover the main three categories of RMs. The first, Scalar RMs, directly output a score (Yuan et al., 2025; Cai et al., 2024; Liu et al., 2024b; Adler et al., 2024; Chen et al., 2025b). The second category, Generative RMs, produces short critiques and may be either strong general-purpose LLMs (Anthropic, 2024; Dubey et al., 2024; Reid et al., 2024; OpenAI, 2023) or fine-tuned GenRMs (Liu et al., 2024b). Lastly, the third type consists of advanced Reasoning RMs based on thinking models (Liu et al., 2025c; Wang et al., 2024; Whitehouse et al., 2025; Chen et al., 2025b; Saha et al., 2025).
Implementation Details. Our implementation leverages a scalable and efficient post-training library, NeMo-RL [^1] https://github.com/NVIDIA-NeMo/RL, with a custom two-turn reward modeling environment. Regarding model design, the penalty constant for format checking is -100 and the weight for reward terms is 10. For training, model-specific learning rates were tuned based on scale: $5\times 10^{-7}$ for Qwen3-8B models (Team, 2025) and $1\times 10^{-6}$ for Qwen3-14B (Team, 2025) models, all trained for 400 optimization steps with validation every 10 steps. For inference, we employ a consistent evaluation protocol with training parameters, computing accuracy as the primary metric, with additional domain-specific metrics for comprehensive analysis. More configurations are in Appendix D.
Table 1: The performance comparison between best-performing baselines. Bold numbers indicate the best performance, Underlined numbers indicate the second best. The more detailed numbers on RewardBench, RM-Bench, and RMB are in Table 6, Table 7, and Table 8 in the Appendix.
Model RewardBench RM-Bench RMB Average
ScalarRMs
SteerLM-RM-70B 88.8 52.5 58.2 66.5
Eurus-RM-7b 82.8 65.9 68.3 72.3
Internlm2-20b-reward 90.2 68.3 62.9 73.6
Skywork-Reward-Gemma-2-27B 93.8 67.3 60.2 73.8
Internlm2-7b-reward 87.6 67.1 67.1 73.9
ArmoRM-Llama3-8B-v0.1 90.4 67.7 64.6 74.2
Nemotron-4-340B-Reward 92.0 69.5 69.9 77.1
Skywork-Reward-Llama-3.1-8B 92.5 70.1 69.3 77.5
INF-ORM-Llama3.1-70B 95.1 70.9 70.5 78.8
GenRMs
Claude-3.5-sonnet-20240620 84.2 61.0 70.6 71.9
Llama3.1-70B-Instruct 84.0 65.5 68.9 72.8
Gemini-1.5-pro 88.2 75.2 56.5 73.3
Skywork-Critic-Llama-3.1-70B 93.3 71.9 65.5 76.9
GPT-4o-0806 86.7 72.5 73.8 77.7
ReasonRMs
JudgeLRM-7B 75.2 64.7 53.1 64.3
DeepSeek-GRM-27B 86.0 72.7 69.0 75.9
Self-taught-evaluator-Llama3.1-70B 90.2 71.4 67.0 76.2
RM-R1-Qwen-Instruct-14B 88.2 76.1 69.2 77.8
RM-R1-DeepSeek-Distilled-Qwen-14B 88.9 81.5 68.5 79.6
EvalPlanner-Llama-3.3-70B-Instruct 93.8 82.1 - -
J1-Llama-70B 93.3 82.7 - -
RM-R1-Qwen-Instruct-32B 91.4 79.1 73.0 81.2
RM-R1-DeepSeek-Distilled-Qwen-32B 90.9 83.9 69.8 81.5
RRM-32B 91.2 84.0 70.0 81.7
Our Methods
BR-RM-Qwen-8B 91.0 85.0 71.8 82.6
BR-RM-Qwen-14B 92.1 85.9 74.7 84.2
4.2 Main Results
Table 1 presents a comprehensive performance comparison of our Branch-and-Rethink Reward Models (BR-RMs) against the best-performing baselines across three standard benchmarks.
ScalarRMs remain competitive but plateau. ScalarRMs such as INF-ORM-Llama3.1-70B achieve the strongest RewardBench score (95.1) and lead most scalar baselines in average accuracy (78.8). However, despite high scores on RewardBench, they lag behind reasoning-focused models on RM-Bench (70.9) and RMB (70.5). Smaller ScalarRMs perform relatively well, suggesting that simply scaling up parameters without reasoning mechanisms brings diminishing returns.
GenRMs improve robustness but trade off consistency. Generative reward models, which critique before scoring, close some gaps but do not consistently surpass scalar baselines. For example, Skywork-Critic-Llama-3.1-70B excels on RewardBench (93.3), yet underperforms on RMB (65.6). GPT-4o achieves balanced results, with the second-best RMB score (73.8) and an average of 77.7, but still lags behind leading scalars and reasoning RMs.
ReasonRMs show clear gains, especially at scale. The RM-R1 model family and related ReasonRMs outperform both ScalarRMs and GenRMs once scaled beyond 14B parameters. RM-R1-DeepSeek-Distilled-Qwen-32B achieves 83.9 on RM-Bench, the strongest baseline prior to our method, with a solid average of 81.5. Similarly, RM-R1-Qwen-Instruct-32B attains 81.2 average. These results confirm that structured reasoning pipelines – when scaled – are effective at overcoming shallow analysis and improving sensitivity to subtle reasoning flaws. However, smaller reasoning RMs (e.g., JudgeLRM-7B at 64.3 average) trail behind even the weaker ScalarRMs, suggesting that naive reasoning without sufficient scale is not sufficient.
BR-RM models consistently perform well across all three benchmarks. While ScalarRMs dominate RewardBench and ReasonRMs lead RM-Bench, our BR-RMs achieve the best overall balance. BR-RM-Qwen-14B reaches 92.1 on RewardBench (close to scalar SOTA), 85.9 on RM-Bench (highest overall), and 74.7 on RMB (top-2 overall), yielding the best average (84.2). Even BR-RM-Qwen-8B, with only 8B parameters, surpasses larger baselines such as GPT-4o and INF-ORM-Llama3.1-70B in average score. These results highlight adaptive two-turn reasoning not only closes the gap with scalar models on factual benchmarks, but also sets a new bar for reasoning robustness.
4.3 Ablation Study
Table 2: Ablation Study. This table summarizes the performance impact of removing key components from our BR-RM models. Values in parentheses show the performance drop from the full models, highlighting the contribution of each component. Due to space constraints, RewardBench is abbreviated as RB.
Prompting strategy BR-RM-Qwen-8B BR-RM-Qwen-14B
RB RMBench RMB Avg RB RMBench RMB Avg
BR-RM (Full Model) 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
Branching Only 90.1 82.6 67.4 80.0 90.9 83.2 71.0 81.7
Unconditioned Rethink 90.6 83.9 69.3 81.3 91.3 84.4 71.9 82.5
Single-Turn Model 90.4 84.5 70.0 81.6 91.3 85.2 73.3 83.3
To validate the contributions of our proposed components, we conduct several ablations based on our BR-RM models. We systematically remove key architectural components to isolate their impact on performance across our evaluation benchmarks (Table 2).
Branching Only In this variant, we remove the second “rethinking” turn. This ablation caused the most significant performance degradation. The first turn is designed for both branching and error detection, which directs its focus toward identifying specific, granular issues. Without the second turn, the model’s final judgment relies entirely on this initial pass. This can mislead the model to over-emphasize details and minor flaws without having a comprehensive view of the whole response. The absence of the holistic re-evaluation step is particularly damaging on nuanced benchmarks like RMB, where the 14B model’s performance drops by a substantial 3.7 points. This shows that the second turn is critical for contextualizing initial findings.
Unconditioned Rethinking Here, we retain the two-turn training but remove branching. The first turn goes through all the dimensions while the second turn is prompted with a generic “re-evaluate the response” instruction. While this model performs better than the “Branching Only” version, it remained well below our full model. For instance, the 14B model sees performance drops of 0.8, 1.5, and 2.8 points on the respective benchmarks. This result underscores the importance of adaptive focus. Without an explicit directive from the first turn, the second pass is less efficient and prone to the same judgment diffusion it is designed to prevent.
Single-Turn Model To test the value of our sequential process, we merge the prompts for both turns into a single instruction. This approach forces the model to rely on only one-time thinking. The resulting prompt is highly complex, demanding that the model handles too many subtasks at the same time: select dimensions, sketch issues, and then re-evaluate to produce a final score. This cognitive overload prevents the model from deeply engaging with any single subtask, leading to a shallower analysis. The consistent, albeit smaller, performance drop confirms that a deliberate, sequential process is critical for high-quality, reasoned evaluation.
4.4 Reward Design Analysis
Table 3: Reward Analysis. Quantitative comparison of our proposed reward design against several ablations. Our approach using a direct binary reward consistently achieves the highest scores. Due to space constraints, RewardBench is abbreviated as RB.
Reward BR-RM-Qwen-8B BR-RM-Qwen-14B
RB RMBench RMB Avg RB RMBench RMB Avg
Our Reward 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
Removing Format Checks 90.5 84.2 71.5 82.1 91.3 85.4 73.1 83.3
Scoring on a Scale 90.0 83.5 70.1 81.2 90.8 85.1 71.7 82.5
Additional Reward for Branch 86.5 79.0 66.2 77.2 87.8 82.7 69.3 79.9
To identify the best training strategy, we evaluated several reward designs in Table 3.
Removing Format Checks First, we tested removing the strict formatting that guides the model’s two-step reasoning process. While this simplified the setup, it allowed the model to find shortcuts, such as producing minimal text just to get the final answer right while skipping the intended reasoning steps. We found that the format check provides crucial guidance and control. It helps the model distinguish between a wrong answer and a badly formatted one; without it, the model would receive the same penalty for incorrect reasoning or just for messy output, which slows down learning. Format checks provide essential structure and unambiguous feedback.
Scoring on a Scale In another experiment, we had the model output a detailed score on a scale, like from -3 to 3. This is inspired by the reward dataset, HelpSteer3, whose human-annotated preferences use a similar scale to express preference strength. During training, the reward was based on the distance between the predicted score and the ground-truth score. To process other datasets with only binary labels, we arbitrarily mapped the winning response to +2 and the losing one to -2. This approach failed, possibly due to a disconnect between the training task and the final evaluation: while training rewarded the model for predicting an exact score, our ultimate evaluation only cared about whether the model correctly identified the winning response (a simple binary choice). This means that the model had no strong incentive to learn the true meaning behind the scores; it only needed to learn if a score was positive or negative. Predicting precise scores is less beneficial when only the winner matters.
Additional Reward for Branching Finally, we tried to reward the first step of the model’s reasoning process individually using a different reward. The key issue here is the absence of a natural reward signal, unlike for the final answer (where ground truth is available). Instead, we calculated it from intermediate values generated during the process. Specifically, we measured the difference of the rewards with and without branching to quantify the effect of Turn 1. But we found that this indirect feedback led to two major practical problems: 1) the approach was computationally expensive, nearly doubling the processing required, and 2) the resulting reward signal was infrequent and erratic, which destabilized training. Calculated intermediate rewards created a noisy, unstable signal.
The Best Reward: Binary Reward with Format Checking. After testing these alternatives, we confirmed that the best reward design is the most straightforward one. In this setup, the model produces a binary choice (A or B) within a required format and receives a simple reward based on whether it is correct. Each of the other designs introduced a flaw – confusing signals, misaligned objectives, or indirect feedback – that hurt performance. This shows that a direct alignment between the task and a reward based on the correct answer provides the most effective foundation for training.
4.5 Training Data Analysis
Table 4: Training Data Analysis. Impact of training data sources for our BR-RM models. Due to space constraints, RewardBench is abbreviated as RB.
Model BR-RM-Qwen-8B BR-RM-Qwen-14B
RB RMBench RMB Avg RB RMBench RMB Avg
Full training data 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
w/o HelpSteer3 90.4 83.5 68.0 80.6 91.5 84.4 70.3 82.1
w/o Skywork 90.6 84.2 66.1 80.3 91.3 85.1 69.7 82.0
w/o Math-Step-DPO 91.0 84.7 70.7 82.1 91.5 85.0 72.6 83.0
w/o Code-Preference 91.1 84.8 71.2 82.4 91.8 85.4 74.0 83.7
To understand each dataset’s contribution, we conducted an ablation study, systematically removing one at a time to measure its impact on model performance. From Table 4, this analysis revealed a clear distinction between the roles of foundational and specialized datasets. For HelpSteer3, removing this foundational dataset led to a broad performance decline, with a 4.4-point drop on the RMB benchmark and a 1.5-point drop on RMBench. This highlights its critical role in establishing general understanding and factual accuracy. For Skywork, excluding our only dataset with safety data caused the most severe impact—a 5.0-point drop on the safety-focused RMB benchmark. This confirms that explicit safety training is essential for model harmlessness. For Math-Step-DPO, removing this specialized dataset predictably weakened reasoning, causing a 2.1-point drop on RMB and a 0.9-point drop on RMBench. This shows in-domain data is necessary for complex, multi-step reasoning. For Code-Preference, the exclusion of this smaller dataset resulted in the most modest decrease (0.7 points on RMB and 0.5 points on RMBench). This suggests its logical reasoning contributions are partially covered by the broader datasets.
5 Conclusion
We recast reward modeling as a focused, two-stage judgment rather than a single holistic score. Our reward model first identifies a small set of instance-critical hypotheses, then performs a targeted reread conditioned on those hypotheses – directing compute where risk is highest. This simple shift curbs judgment diffusion, reduces style-driven bias, and exposes subtle factual or logical slips, yielding more stable preferences across benchmarks. In essence, the same “think twice” discipline that improves solvers also makes their judges more reliable. Extensions of this work may include integrating verification tools like retrieval and code runners to check claims, automatically generating dynamic evaluation rules to replace static rubrics, and using model uncertainty to adaptively decide when and how deeply to re-evaluate an output.
Appendix A Preliminary Study on Judgment Diffusion
Figure 3: Comparing token allocation between our method and a recent ReasonRM (13) on four subsets from RM-Bench. Our method adaptively focuses its generative analysis on a few critical dimensions for each task (e.g., Information Accuracy for chat), while the baseline spreads its tokens broadly across many criteria.
We design a preliminary study to quantify the focus of RMs on multiple evaluation dimensions by analyzing the token allocation of model generation to verify our concepts of focus dilution and shallow analysis.
Our experimental setup involves comparing two models head-to-head. The first, our baseline, RRM-32B (13), is a standard generative RM trained to produce a holistic critique in a single forward pass, thereby considering all evaluation criteria at once. The second is our proposed two-turn branch-and-rethink approach, which executes both the Adaptive Branching and Branch-Conditioned Rethinking steps. Crucially, both models are given the exact same maximum token limit, 16384. This ensures we’re measuring how they choose to spend a fixed test-time compute budget. For the dataset, we selected all the subsets from the RM-Bench benchmark to test distinct capabilities. From the Chat split, we sample 100 instances manually that require sharp factual accuracy and solid logical reasoning to judge correctly. From the Code split, we sampled 100 instances where the judgment hinges on catching subtle bugs, checking algorithmic correctness, and validating the implementation. From the Chat split, we sample 100 general conversational instances that demand strong factuality and response completeness. From the Math split, we sample 100 mathematical reasoning problems where computational precision and logical correctness are paramount. From the Safety split, we sampled 100 instances where identifying potential harms, ethical violations, and safety-critical issues is paramount.
We follow a clear procedure to gather our results. First, for every task in our test sets, we have both the Baseline and Our Method generate their respective judgments. The next and most critical step is Token Allocation. Here, we prompt a large language model, GPT-4o-0806, to read the generated text from each model and attribute every single sentence to one of the predefined evaluation dimensions (e.g., Information Accuracy, Logical Reasoning). For our method, we combine the tokens from both the initial Adaptive Branching turn and the final Branch-Conditioned Rethinking turn. After that, we calculate the proportions by determining the percentage of total test-time compute (tokens) spent on each dimension for every instance. Finally, we average these percentage allocations across all instances within each split.
As shown in Figure 3, the baseline model exemplifies the problems of judgment diffusion. For example, for the (a) Chat and (b) Code tasks, it distributes its tokens broadly across numerous dimensions, such as Writing Clarity and Content Relevance, even when they are not critical considering the strong power of current LLMs. This is a clear indicator of focus dilution, as the model’s limited generative budget is spread thin instead of being concentrated on what matters. This, in turn, leads to a shallow analysis of the most important factors. For instance, in the Code task, the baseline allocates only about 25% of its tokens to Implementation Capability. In contrast, our method demonstrates adaptive focus: it concentrates over 70% of its generation on Information Accuracy and Logical Reasoning for Chat and over 75% on Implementation Capability and Computational Precision for Code. This evidence suggests that without a structured mechanism for focusing, RMs default to a diffuse and superficial evaluation, whereas our approach enables a targeted and deeper analysis.
Appendix B Experiment Setting
B.1 Datasets
We train on a diverse mixture of datasets to ensure both coverage and domain specialization:
  • HelpSteer3 (42): A large-scale, human-annotated dataset providing diverse open-domain preference pairs across multiple languages and tasks. Each pair is supplemented with fine-grained ranking scores (from -3 to 3). This dataset contributes the breadth needed to capture general human preferences in everyday conversational and instructional settings.
  • Skywork Reward Preference-80K (19): A broad-coverage dataset emphasizing instruction-following, reasoning quality, and safety compliance. Unlike HelpSteer3, which covers open-ended domains, Skywork’s annotations focus on alignment with user intent and responsible behavior. This introduces strong supervision for safety- and adherence-related dimensions, which are often underrepresented in general instruction datasets.
  • Code-Preference-Pairs (8K) (39): A curated collection of programming-related preference pairs, highlighting tradeoffs between functional correctness, readability, and efficiency. This dataset provides focused supervision for Implementation Capability and Computational Precision, two dimensions frequently selected in adaptive branching for code tasks. Without such domain-specific data, reward models tend to overfit to stylistic features instead of true semantic correctness.
  • Math-Step-DPO-10K (15): Preference pairs drawn from mathematical problem solving, where solutions may differ in intermediate reasoning steps as well as final correctness. This dataset is critical for training models to detect subtle reasoning errors, making it an ideal testbed for Logical Reasoning and Information Accuracy. It encourages the model not only to verify final answers but also to scrutinize the validity of intermediate steps.
B.2 Evaluation Data
Following Chen et al. (9), we evaluate our models on three widely used and challenging benchmarks that capture different facets of alignment quality. Together, they provide a comprehensive testbed spanning subtle reasoning, real-world coverage, and robustness to bias.
  • RewardBench (17): A standard benchmark for evaluating reward models across diverse domains, including factual QA, multi-step reasoning, safety-critical instructions, and stylistic tradeoffs. It is widely adopted for measuring whether reward models align with human preferences in broad settings. RewardBench serves as our measure of generalization breadth: models must balance correctness, reasoning, safety, and fluency simultaneously.
  • RM-Bench (21): A more fine-grained evaluation that emphasizes sensitivity to subtle errors such as quiet factual slips, brittle logical steps, and superficial stylistic cues. RM-Bench is also designed to probe systematic biases, including verbosity bias (favoring longer responses) and position bias (favoring earlier/later candidates). Performance here reflects a model’s ability to avoid shallow scanning and resist bias—precisely where our two-turn “branch-and-rethink” framework aims to excel.
  • RMB (49): A large-scale, comprehensive benchmark covering 49 real-world scenarios, ranging from open-domain dialogue and question answering to code, math, and safety. Unlike other benchmarks, RMB supports both pairwise preference comparisons and Best-of-N (BoN) evaluation, where models must select the best among multiple candidates. This setup better reflects practical deployment conditions, since real-world LLMs often generate diverse candidate sets. RMB explicitly reports performance on helpfulness and harmlessness, making it a balanced measure of alignment quality across both capability and safety.
B.3 Baselines
We compare against a wide array of state-of-the-art reward models, grouped by their architecture and output modality. This list is largely inspired by the closely related prior work on RM-R1 (9):
  • Scalar RMs: Conventional reward models that directly output a single scalar score without generating intermediate reasoning. We compare to SteerLM-RM-70B, Eurus-RM-7B (47), InternLM2-20B-Reward, InternLM2-7B-Reward (6), Skywork-Reward-Gemma-2-27B , Skywork-Reward-Llama-3.1-8B (19), ArmoRM-Llama3-8B-v0.1, Nemotron-4-340B-Reward (1), and INFORM-Llama3.1-70B (9).
  • Generative RMs (GenRMs): Reward models that produce natural language critiques or rationales before collapsing them into a scalar score. This category includes both closed-source and open-source representatives: Claude-3.5-Sonnet (3), Llama-3.1-70B-Instruct (12), Gemini-1.5-Pro (28), GPT-4o (25), and Skywork-Critic-Llama-3.1-70B (19).
  • Reasoning RMs (ReasonRMs): Advanced reward models that explicitly perform intermediate reasoning or structured critique before producing a preference. This includes JudgeLRM (8), DeepSeek-PairRM-27B, DeepSeek-GRM-27B-RFT, DeepSeek-GRM-27B (22), Self-Taught-Evaluator-Llama3.1-70B (40), J1-Llama-70B (44), EvalPlanner-Llama-3.3-70B-Instruct (29), as well as the RM-R1 model family (9).
We prioritize official results reported by model providers or other published works, including (9) and (20), and evaluate remaining models under their standardized setup (including RRM-32B and Self-taught-evaluator-Llama3.1-70B).
Appendix C Prompts Used in BR-RM
[…]
Figure 4: Prompt for adaptive branching.
Figure 5: Prompt for branch-conditioned rethinking.
Appendix D Implementation Details
Table 5: Training and Inference Configurations.
Category Parameter Value
Distributed Training Total Nodes 8
GPUs per Node 8
Total GPUs 64
Tensor Parallelism 8
Context Parallelism 1
Pipeline Parallelism 1
Memory Optimization FSDP2 (DTensor)
Training Hyperparameters Optimization Algorithm GRPO
Training Steps 400
Global Batch Size 256
Micro Batch Size 1
Prompts per Step 128
Generations per Prompt 8
Temperature 1.0
PPO Clip Min 0.2
PPO Clip Max 0.28
KL Penalty 0.001
Weight Decay 0.0
Gradient Clip Norm 1.0
Adam $\beta$ 0.9
Adam $\epsilon$ $1\times 10^{-8}$
Learning Rate (8B) $5\times 10^{-7}$
Learning Rate (14B) $1\times 10^{-6}$
Max Total Tokens 16384
Format Penalty 100-100
Reward Weight 10
Precision bfloat16
Evaluation Hyperparameters Batch Size 1,024
Generation Settings Backend vLLM
Top-p 0.95
Top-k 20
Temperature 1.0
Max New Tokens 8,192
Max Total Tokens 16,384 tokens
Our implementation leverages a scalable and efficient post-training library, NeMo-RL [^2] https://github.com/NVIDIA-NeMo/RL, with a custom two-turn reward modeling environment. Regarding model design, the penalty constant for format checking is -100. The prompts used are in Appendix C. The training pipeline processes each stage sequentially, applying stage-specific stop strings and parsing structured outputs using regex-based extraction. We employ Group Relative Policy Optimization (GRPO, Shao et al. 30), which extends PPO by optimizing over groups of responses simultaneously, providing more stable gradients for preference learning. The distributed training configuration utilizes 8 nodes with 8 GPUs each, employing fully sharded data parallel with tensor parallelism of 8 across GPUs. Memory optimization techniques include CPU offloading for optimizer states, activation checkpointing, and mixed precision training with bfloat16. The training process uses asymmetric GRPO clipping bounds [0.2, 0.28] for variance reduction and a KL divergence penalty of 0.001. Model-specific learning rates were tuned based on scale: $5\times 10^{-7}$ for 8B models and $1\times 10^{-6}$ for 14B models, all trained for 400 optimization steps with validation every 5 steps. For inference, we employ vLLM with async engine support for high-throughput evaluation. The evaluation protocol maintains consistency with training parameters, using temperature 1.0 for generation and computing accuracy as the primary metric, with additional domain-specific metrics for comprehensive analysis. More configuration details are shown in Table 5 in the Appendix. Table 5 presents the key hyperparameters used for training our models. These parameters were carefully selected to optimize the reinforcement learning process and ensure effective development of reasoning capabilities in our models.
Appendix E Detailed Experiment Results
Table 6: Results on RewardBench.
Model Chat Chat_Hard Safety Reasoning Overall
ScalarRMs
Eurus-RM-7b 98.0 65.6 81.4 86.3 82.8
Internlm2-7b-reward 99.2 69.5 87.2 94.5 87.6
SteerLM-RM-70B 91.3 80.3 92.8 90.6 88.8
Cohere-0514 96.4 71.3 92.3 97.7 89.4
Internlm2-20b-reward 98.9 76.5 89.5 95.8 90.2
ArmorRM-Llama3-8B-v0.1 96.9 76.8 90.5 97.3 90.4
Nemotron-4-340B-Reward 95.8 87.1 91.5 93.6 92.0
Skywork-Reward-Llama-3.1-8B 95.8 87.3 90.8 96.2 92.5
Skywork-Reward-Gemma-2-27B 95.8 91.4 91.9 96.1 93.8
INF-ORM-Llama3.1-70B 96.6 91.0 93.6 99.1 95.1
GenRMs
Llama3.1-8B-Instruct 85.5 48.5 75.6 72.1 70.4
Prometheus-7B-v2 93.0 47.1 80.5 77.4 74.5
Llama3.1-70B-Instruct 97.2 70.2 82.8 86.0 84.0
Llama3.1-405B-Instruct 97.2 74.6 77.6 87.1 84.1
Claude-3-5-sonnet-20240620 96.4 74.0 81.6 84.7 84.2
GPT-4o-0806 96.1 76.1 86.6 88.1 86.7
Gemini-1.5-pro 92.3 80.6 87.9 92.0 88.2
SFR-LlaMa-3.1-70B-Judge-r 96.9 84.8 91.6 97.6 92.7
Skywork-Critic-Llama-3.1-70B 96.6 87.9 93.1 95.5 93.3
ReasonRMs
JudgeLRM 92.9 56.4 78.2 73.6 75.2
SynRM 38.0 82.5 74.1 87.1 70.4
RM-R1-DeepSeek-DISTILLED-Qwen-7B 88.9 66.2 78.4 87.0 80.1
Cloud 97.0 58.0 84.0 92.0 82.8
DeepSeek-GRM-16B 90.8 74.3 84.7 81.8 82.9
DeepSeek-GRM-27B-RFT 94.7 77.2 87.0 79.2 84.5
RM-R1-Qwen-INSTRUCT-7B 94.1 74.6 85.2 86.7 85.2
DeepSeek-GRM-27B 94.1 78.3 88.0 83.8 86.0
DeepSeek-PairRM-27B 95.3 86.8 52.3 92.0 87.1
RM-R1-Qwen-INSTRUCT-14B 93.6 80.5 86.9 92.0 88.2
RM-R1-DeepSeek-DISTILLED-Qwen-14B 91.3 79.4 89.3 95.5 88.9
Self-taught-evaluator-Llama3.1-70B 96.9 83.1 89.6 88.4 90.0
RM-R1-DeepSeek-DISTILLED-Qwen-32B 95.3 80.3 91.1 96.8 90.9
RRM-32B 94.7 81.1 90.7 98.3 91.2
RM-R1-Qwen-INSTRUCT-32B 95.3 83.1 91.9 95.2 91.4
J1-Llama-70B 96.1 90.1 91.9 94.9 93.3
EvalPlanner-Llama-3.3-70B-Instruct 97.7 89.5 91.7 96.1 93.8
Our Models
BR-RM-Qwen-8B 95.8 80.1 90.4 97.5 91.0
BR-RM-Qwen-14B 97.0 82.4 90.0 98.8 92.1
In this section, we provide the full experiment results and a more comprehensive coverage of existing baselines. The results of RewardBench, RM-Bench, and RMB are provided in Tables 6, 7, and 8.
Table 7: The full results of tested reward models on RM-Bench. Chat, Math, Code, and Safety show the model’s Average accuracy on each domain. Easy, Normal, Hard show the model’s Accuracy on each difficulty level across all domains. Bold numbers indicate the best performance, Underlined numbers indicate the second best.
Model Chat Math Code Safety Easy Normal Hard Avg
ScalarRMs
steerlm-70b 56.4 53.0 49.3 51.2 48.3 54.9 54.3 52.5
tulu-v2.5-70b-preference-mix-rm 58.2 51.4 55.5 87.1 72.8 65.6 50.7 63.0
Mistral-7B-instruct-Unified-Feedback 56.5 58.0 51.7 86.8 87.1 67.3 35.3 63.2
RM-Mistral-7B 57.4 57.0 52.7 87.2 88.6 67.1 34.9 63.5
Eurus-RM-7b 59.9 60.2 56.9 86.5 87.2 70.2 40.2 65.9
internlm2-7b-reward 61.7 71.4 49.7 85.5 85.4 70.7 45.1 67.1
Skywork-Reward-Gemma-2-27B 69.5 54.7 53.2 91.9 78.0 69.2 54.9 67.3
ArmorRM-Llama3-8B-v0.1 67.8 57.5 53.1 92.4 82.2 71.0 49.8 67.7
GRM-llama3-8B-sftreg 62.7 62.5 57.8 90.0 83.5 72.7 48.6 68.2
internlm2-20b-reward 63.1 66.8 56.7 86.5 82.6 71.6 50.7 68.3
llama-3-OffsetBias-RM-8B 71.3 61.9 53.2 89.6 84.6 72.2 50.2 69.0
Nemotron-340B-Reward 71.2 59.8 59.4 87.5 81.0 71.4 56.1 69.5
URM-LLaMA-3.1-8B 71.2 61.8 59.1 93.1 84.0 73.2 53.0 70.0
Skywork-Reward-Llama-3.1-8B 69.5 60.6 54.5 95.7 89.0 74.7 46.6 70.1
INF-ORM-Llama3.1-70B 66.3 65.6 56.8 94.8 91.8 76.1 44.8 70.9
GenRMs
tulu-v2.5-dpo-13b-chatbot-arena-2023 64.9 52.3 50.5 62.3 82.8 60.2 29.5 57.5
tulu-v2.5-dpo-13b-nectar-60k 56.3 52.4 52.6 73.8 86.7 64.3 25.4 58.8
stablelm-2-12b-chat 67.2 54.9 51.6 65.2 69.1 63.5 46.6 59.7
tulu-v2.5-dpo-13b-stackexchange-60k 66.4 49.9 54.2 69.0 79.5 63.0 37.2 59.9
Nous-Hermes-2-Mistral-7B-DPO 58.8 55.6 51.3 73.9 69.5 61.1 49.1 59.9
Claude-3-5-sonnet-20240620 62.5 62.6 54.4 64.4 73.8 63.4 45.9 61.0
tulu-v2.5-dpo-13b-$h$-rlhf-60k 68.4 51.1 52.3 76.5 53.6 63.0 69.6 62.1
tulu-2-dpo-13b 66.4 51.4 57.8 85.4 86.9 66.7 37.7 63.8
SOLAR-10.7B-Instruct-v1.0 78.6 52.3 49.6 78.9 57.5 67.6 69.4 64.8
Llama3.1-70B-Instruct 64.3 67.3 47.3 83.0 74.7 67.8 54.1 65.5
Skywork-Critic-Llama-3.1-70B 71.4 64.0 56.8 94.8 85.6 73.7 56.5 71.9
GPT-4o-0806 67.2 67.5 63.6 91.7 83.4 75.6 58.7 72.5
Gemini-1.5-pro 71.6 73.9 63.7 91.3 83.1 77.6 64.7 75.2
ReasonRMs
DeepSeek-GRM-27B 72.4 59.2 68.0 91.1 84.6 76.5 57.0 64.7
JudgeLRM 59.9 59.9 51.9 87.3 73.2 66.2 54.8 64.7
RM-R1-Qwen-INSTRUCT-7B 66.6 67.0 54.6 92.6 79.2 71.7 59.7 70.2
Self-taught-evaluator-llama3.1-70B 73.4 65.7 56.3 90.4 80.2 74.5 59.7 71.5
RM-R1-DeepSeek-DISTILLED-Qwen-7B 64.0 83.9 56.2 85.3 75.9 73.1 68.1 72.4
RM-R1-Qwen-INSTRUCT-14B 75.6 75.4 60.6 93.6 82.6 77.5 68.8 76.1
RM-R1-Qwen-INSTRUCT-32B 75.3 80.2 66.8 93.9 86.3 80.5 70.4 79.1
RM-R1-DeepSeek-DISTILLED-Qwen-32B 71.8 90.5 69.5 94.1 86.2 83.6 74.4 81.5
EvalPlanner-Llama-3.3-70B-Instruct 81.1 80.8 84.3 82.1
RM-R1-DeepSeek-DISTILLED-Qwen-14B 74.2 91.8 74.1 95.4 89.5 85.4 76.7 83.9
RRM-32B 73.7 91.9 75.0 95.3 90.7 85.3 76.3 84.0
Our Models
BR-RM-Qwen-8B 76.4 94.1 77.0 92.7 91.7 87.3 76.1 85.0
BR-RM-Qwen-14B 77.3 92.6 79.8 93.7 92.0 88.1 77.6 86.1
Table 8: The leaderboard of RMB, ranked by the average score of all subsets. Bold numbers indicate the best performance, Underlined numbers indicate the second best.
Model Helpfulness Harmlessness
BoN Pairwise BoN Pairwise Overall
ScalarRMs
Tulu-v2.5-13b-preference-mix-rm 0.355 0.562 0.351 0.545 0.453
SteerLM-RM 70B 0.502 0.574 0.578 0.673 0.582
Skywork-Reward-Gemma-2-27B 0.472 0.653 0.561 0.721 0.602
Internlm2-20b-reward 0.585 0.763 0.499 0.670 0.629
ArmorRM-Llama3-8B-v0.1 0.636 0.787 0.497 0.663 0.646
Internlm2-7b-reward 0.626 0.782 0.563 0.712 0.671
Eurus-RM-7b 0.679 0.818 0.543 0.693 0.683
Skywork-Reward-Llama-3.1-8B 0.627 0.781 0.603 0.759 0.693
INF-ORM-Llama3.1-70B 0.650 0.798 0.607 0.767 0.705
Starling-RM-34B 0.604 0.774 0.674 0.795 0.712
GenRMs
Llama2-70b-chat 0.289 0.613 0.249 0.602 0.438
Llama3.1-8B-Instruct 0.365 0.675 0.267 0.653 0.490
Gemini-1.5-pro 0.536 0.763 0.299 0.661 0.565
Mixtral-8x7B-Instruct-v0.1 0.480 0.706 0.491 0.671 0.587
skywork-critic-llama3.1-8B 0.600 0.725 0.578 0.578 0.620
skywork-critic-llama3.1-70B 0.640 0.753 0.614 0.614 0.655
Llama3.1-70B-Instruct 0.648 0.811 0.558 0.739 0.689
Mistral-Large-2407 0.678 0.817 0.583 0.725 0.701
Claude-3-5-sonnet 0.705 0.838 0.518 0.764 0.706
Qwen2-72B-Instruct 0.645 0.810 0.649 0.789 0.723
GPT-4o-2024-05-13 0.639 0.815 0.682 0.814 0.738
ReasonRMs
JudgeLRM 0.363 0.699 0.363 0.674 0.531
RM-R1-DeepSeek-DISTILLED-Qwen-7B 0.451 0.658 0.429 0.664 0.551
RM-R1-Qwen-INSTRUCT-7B 0.543 0.740 0.608 0.765 0.664
Self-taught-evaluator-llama3.1-70B 0.616 0.786 0.546 0.733 0.670
Deepseek-27B-RFT 0.592 0.801 0.548 0.765 0.670
RM-R1-DeepSeek-DISTILLED-Qwen-14B 0.593 0.765 0.613 0.769 0.685
Deepseek-GRM-27B 0.623 0.805 0.570 0.761 0.690
RM-R1-Qwen-INSTRUCT-14B 0.594 0.776 0.620 0.778 0.692
RM-R1-DeepSeek-DISTILLED-Qwen-32B 0.620 0.782 0.618 0.771 0.698
RRM-32B 0.609 0.763 0.661 0.766 0.700
RM-R1-Qwen-INSTRUCT-32B 0.636 0.791 0.682 0.809 0.730
Our Models
BR-RM-Qwen-8B 0.597 0.769 0.685 0.822 0.718
BR-RM-Qwen-14B 0.670 0.810 0.693 0.816 0.747
Appendix F Case Study
To demonstrate the effectiveness of our Branch-and-Rethink Reasoning Reward Model, we present a case study involving the evaluation of C++ function implementations from RM-Bench. The user query requests a function unique_digits(vector<int> x) that returns a sorted vector containing only integers without any even digits. This programming task represents a typical scenario where multiple correct implementations may exist with subtle quality differences that traditional reward models might overlook.
Turn 1: Adaptive Branching
In the first turn, our model performs adaptive branching by dynamically selecting the most relevant evaluation dimensions for this specific task. Rather than applying a fixed set of criteria, the model identifies three critical dimensions based on the task context:
  1. 1.
    Implementation Capability – Assessing the correctness of the algorithmic logic and code structure
  2. 2.
    Computational Precision – Evaluating the accuracy of digit-checking operations and numerical computations
  3. 3.
    Instruction Adherence – Verifying compliance with specified requirements, including sorting and filtering constraints
During this phase, the model conducts focused analysis on both responses. For Response 1, the model identifies a properly structured helper function for digit checking, correct filtering logic, and the inclusion of a complete main function with example usage. Response 2 exhibits functionally identical core logic but presents a commented-out main function. Notably, both responses pass the initial screening with no critical errors identified, a scenario that challenges traditional single-pass evaluation methods.
Turn 2: Branch-Conditioned Rethinking
The second turn leverages the insights from adaptive branching to conduct deeper, issue-driven analysis. Based on the absence of critical errors identified in Turn 1, the model applies a hierarchical evaluation framework prioritizing Correctness >> Process >> Presentation. This accuracy-critical hierarchy ensures that functional correctness remains paramount while still distinguishing between responses of similar quality.
The branch-conditioned rethinking reveals subtle but meaningful differences:
  • Response 1 provides a fully functional implementation with an active main function and verifiable example cases, demonstrating completeness in both solution and presentation
  • Response 2 offers equivalent algorithmic correctness but with marginally reduced completeness due to the commented-out main function, which may impact immediate usability
Through this two-turn process, the model generates a nuanced ranking (Response 1 is better) that captures quality distinctions beyond binary correctness. This granular evaluation is crucial for reinforcement learning applications where the reward signal must differentiate between multiple acceptable solutions to guide model improvement effectively.
Figure 6: Case Study on C++ Function Implementations. The framework operates in two sequential turns: (a) Turn 1: Adaptive Branching, where the model dynamically selects critical evaluation dimensions (Implementation Capability, Computational Precision, and Instruction Adherence) based on the user query context and performs focused analysis to detect specific issues; (b) Turn 2: Branch-Conditioned Rethinking, where the model conducts deeper, issue-driven re-evaluation using a hierarchical assessment framework (Correctness >> Process >> Presentation) to generate final ranking scores. The case study demonstrates the evaluation of two functionally similar responses to a unique_digits() implementation task, showing how the model’s two-turn approach enables nuanced quality assessment despite both responses having no critical errors. The final reward judgment (Response 1 ranked higher) is determined by subtle presentation and completeness differences identified through the branch-conditioned rethinking process, illustrating the model’s capability to provide fine-grained distinctions necessary for effective reinforcement learning optimization.
Appendix G Statement on Usage of Large Language Models
In the preparation of this submission, large language models were utilized as a writing assistance tool. The primary role of the LLM was to enhance the clarity, grammar, and readability of the text. Its use was confined to language polishing, such as correcting syntax and improving sentence structure. The LLM was not used for any core research activities, including ideation, data analysis, or the formulation of conclusions. All intellectual contributions, including the research questions, methodology, and scientific insights, are the original work of the human authors. We have meticulously reviewed and edited all text and assume full responsibility for the final content of this paper.

### v3

@@
 Experimental results demonstrate that our model achieves state-of-the-art performance on three challenging reward modeling benchmarks across diverse domains.
-The code and the models will be released soon.
 Source Code: https://github.com/yzjiao/BR-RM.
@@
 Recent “thinking” models and analyses of test-time compute further highlight the benefits of deliberate, multi-step inference and adaptive allocation (Snell et al., 2024).
+Figure 1: Comparing token allocation between our method and a recent ReasonRM (Guo et al., 2025) on two subsets from RM-Bench.
+Our method adaptively focuses its generative analysis on a few critical dimensions for each task (e.g., Information Accuracy for chat), while the baseline spreads its tokens broadly across many criteria.
+More details in Appendix 7.
 Yet a single explicit-reasoning pass is often brittle – a quiet factual slip, a local contradiction, or a small logic bug can persist because attention is spread across many potential concerns.
@@
 Thus, even as they add reasoning, they do not guarantee that attention concentrates where risk is highest.
-Figure 1: Comparing token allocation between our method and a recent ReasonRM (Guo et al., 2025) on two subsets from RM-Bench.
-Our method adaptively focuses its generative analysis on a few critical dimensions for each task (e.g., Information Accuracy for chat), while the baseline spreads its tokens broadly across many criteria.
-More details in Appendix 3.
 We highlight a recurring weakness across existing RMs, which we call judgment diffusion.
@@
 Extensive experiments on reward modeling benchmarks show that our method, BR-RM, achieves overall state-of-the-art performance across RewardBench (Lambert et al., 2025b), RMBench (Liu et al., 2025b), and RMB (Zhou et al., 2025), outperforming strong baselines across multiple domains, including reasoning, general knowledge, safety, and alignment with human preferences.
-Beyond final performance, we conduct extensive empirical analyses of BR-RM, including the ablation study of our training recipes, the detailed analysis of the reward designs, and concrete impacts of training data sources.
+Beyond final performance, we conduct extensive empirical analyses, including the ablation study of our training recipes, the detailed analysis of the reward designs, and concrete impacts of training data sources.
 These insights provide a deeper understanding of reward reasoning processes and will hopefully inspire the development of future reasoning reward models.
@@
 This Branch-Conditioned Rethinking is not an independent assessment but a deliberate critique of the responses through the lens of the criteria identified in Stage 1.
-To ensure this deliberation is principled, we incorporate task-specific evaluation hierarchies, $\mathcal{H}_{\text{task}}$, which provide a structured inductive bias.
-These hierarchies encode a stable ordering of importance for different criteria, such as:
-  • Accuracy-Critical Tasks: $\text{Correctness}>\text{Methodology}>\text{Clarity}$
-  • Creative Tasks: $\text{Intent Alignment}>\text{Quality}>\text{Novelty}$
+To ensure this deliberation is principled, we incorporate task-specific evaluation hierarchies $\mathcal{H}_{\text{task}}$ to provide a structured inductive bias.
+For example, accuracy-driven tasks prioritize Correctness and Methodology over Clarity, whereas creative tasks favor Intent Alignment and Novelty.
 The model is conditioned on $\tau_{1}$ and guided by $\mathcal{H}_{\text{task}}$ to generate the second part of the trace, $\tau_{2}$, which contains a detailed comparative judgment and the final decision, $\hat{z}$.
@@
 We employ Generalized Reward Policy Optimization (GRPO, Shao et al.
-2024), a variant of PPO, chosen for its training stability and sample efficiency.
+2024), a variant of PPO, chosen for its training stability and efficiency.
 The training loop (Algorithm 1) optimizes the policy $\pi_{\theta}$ to maximize a composite reward function over the distribution of generated traces.
 The GRPO objective balances reward maximization with training stability via a clipped surrogate objective and a KL-divergence penalty against a reference policy $\pi_{\text{ref}}$:
-$L_{\text{GRPO}}=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\min\left(\rho(\tau)A(\tau),\text{clip}(\rho(\tau),1-\epsilon,1+\epsilon)A(\tau)\right)\right]-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$ (3)
+$L_{\text{GRPO}}={}\mathbb{E}_{\tau\sim\pi_{\theta}}\Big[\min\Big(\rho(\tau)A(\tau),$ (3)
+$\quad\text{clip}(\rho(\tau),1-\epsilon,1+\epsilon)A(\tau)\Big)\Big]-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$
 where $\rho(\tau)=\frac{\pi_{\theta}(\tau)}{\pi_{\text{ref}}(\tau)}$ is the probability ratio.
@@
 Regarding model design, the penalty constant for format checking is -100 and the weight for reward terms is 10.
-For training, model-specific learning rates were tuned based on scale: $5\times 10^{-7}$ for Qwen3-8B models (Team, 2025) and $1\times 10^{-6}$ for Qwen3-14B (Team, 2025) models, all trained for 400 optimization steps with validation every 10 steps.
+For training, model-specific learning rates were tuned based on scale: $5\times 10^{-7}$ for Qwen3-8B models (Team, 2025) and $1\times 10^{-6}$ for Qwen3-14B (Team, 2025) models, all trained for 400 optimization steps.
 For inference, we employ a consistent evaluation protocol with training parameters, computing accuracy as the primary metric, with additional domain-specific metrics for comprehensive analysis.
 More configurations are in Appendix D.
-Table 1: The performance comparison between best-performing baselines.
-Bold numbers indicate the best performance, Underlined numbers indicate the second best.
-The more detailed numbers on RewardBench, RM-Bench, and RMB are in Table 6, Table 7, and Table 8 in the Appendix.
+Table 1: The performance comparison between best-performing baselines. Bold and underline indicate the first and second best performance respectively.
 Model RewardBench RM-Bench RMB Average
 Scalar RMs
-SteerLM-RM-70B 88.8 52.5 58.2 66.5
-Eurus-RM-7b 82.8 65.9 68.3 72.3
-Internlm2-20b-reward 90.2 68.3 62.9 73.6
-Skywork-Reward-Gemma-2-27B 93.8 67.3 60.2 73.8
-Internlm2-7b-reward 87.6 67.1 67.1 73.9
-ArmoRM-Llama3-8B-v0.1 90.4 67.7 64.6 74.2
-Nemotron-4-340B-Reward 92.0 69.5 69.9 77.1
-Skywork-Reward-Llama-3.1-8B 92.5 70.1 69.3 77.5
-INF-ORM-Llama3.1-70B 95.1 70.9 70.5 78.8
+SteerLM-RM-70B (Chen et al., 2025b) 88.8 52.5 58.2 66.5
+Eurus-RM-7b (Yuan et al., 2025) 82.8 65.9 68.3 72.3
+Internlm2-20b-reward (Cai et al., 2024) 90.2 68.3 62.9 73.6
+Skywork-Reward-Gemma-2-27B (Liu et al., 2024b) 93.8 67.3 60.2 73.8
+Internlm2-7b-reward (Cai et al., 2024) 87.6 67.1 67.1 73.9
+ArmoRM-Llama3-8B-v0.1 (Chen et al., 2025b) 90.4 67.7 64.6 74.2
+Nemotron-4-340B-Reward (Adler et al., 2024) 92.0 69.5 69.9 77.1
+Skywork-Reward-Llama-3.1-8B (Liu et al., 2024b) 92.5 70.1 69.3 77.5
+INF-ORM-Llama3.1-70B (Chen et al., 2025b) 95.1 70.9 70.5 78.8
 Gen RMs
-Claude-3.5-sonnet-20240620 84.2 61.0 70.6 71.9
-Llama3.1-70B-Instruct 84.0 65.5 68.9 72.8
-Gemini-1.5-pro 88.2 75.2 56.5 73.3
-Skywork-Critic-Llama-3.1-70B 93.3 71.9 65.5 76.9
-GPT-4o-0806 86.7 72.5 73.8 77.7
+Claude-3.5-sonnet (Anthropic, 2024) 84.2 61.0 70.6 71.9
+Llama3.1-70B-Instruct (Dubey et al., 2024) 84.0 65.5 68.9 72.8
+Gemini-1.5-pro (Reid et al., 2024) 88.2 75.2 56.5 73.3
+Skywork-Critic-Llama-3.1-70B (Liu et al., 2024b) 93.3 71.9 65.5 76.9
+GPT-4o-0806 (OpenAI, 2023) 86.7 72.5 73.8 77.7
 Reason RMs
-JudgeLRM-7B 75.2 64.7 53.1 64.3
-DeepSeek-GRM-27B 86.0 72.7 69.0 75.9
-Self-taught-evaluator-Llama3.1-70B 90.2 71.4 67.0 76.2
-RM-R1-Qwen-Instruct-14B 88.2 76.1 69.2 77.8
-RM-R1-DeepSeek-Distilled-Qwen-14B 88.9 81.5 68.5 79.6
-EvalPlanner-Llama-3.3-70B-Instruct 93.8 82.1 - -
-J1-Llama-70B 93.3 82.7 - -
-RM-R1-Qwen-Instruct-32B 91.4 79.1 73.0 81.2
-RM-R1-DeepSeek-Distilled-Qwen-32B 90.9 83.9 69.8 81.5
-RRM-32B 91.2 84.0 70.0 81.7
+JudgeLRM-7B (Chen et al., 2025a) 75.2 64.7 53.1 64.3
+DeepSeek-GRM-27B (Liu et al., 2025c) 86.0 72.7 69.0 75.9
+Self-taught-evaluator-70B (Wang et al., 2024) 90.2 71.4 67.0 76.2
+RM-R1-Qwen-14B (Chen et al., 2025b) 88.2 76.1 69.2 77.8
+RM-R1-DS-Distill-14B (Chen et al., 2025b) 88.9 81.5 68.5 79.6
+EvalPlanner-Llama-3.3-70B (Saha et al., 2025) 93.8 82.1
+J1-Llama-70B (Whitehouse et al., 2025) 93.3 82.7
+RM-R1-Qwen-32B (Chen et al., 2025b) 91.4 79.1 73.0 81.2
+RM-R1-DS-Distill-32B (Chen et al., 2025b) 90.9 83.9 69.8 81.5
+RRM-32B (Guo et al., 2025) 91.2 84.0 70.0 81.7
 Our Methods
@@
 4.2 Main Results
-Table 1 presents a comprehensive performance comparison of our Branch-and-Rethink Reward Models (BR-RMs) against the best-performing baselines across three standard benchmarks.
-ScalarRMs remain competitive but plateau.
-ScalarRMs such as INF-ORM-Llama3.1-70B achieve the strongest RewardBench score (95.1) and lead most scalar baselines in average accuracy (78.8).
-However, despite high scores on RewardBench, they lag behind reasoning-focused models on RM-Bench (70.9) and RMB (70.5).
-Smaller ScalarRMs perform relatively well, suggesting that simply scaling up parameters without reasoning mechanisms brings diminishing returns.
-GenRMs improve robustness but trade off consistency.
-Generative reward models, which critique before scoring, close some gaps but do not consistently surpass scalar baselines.
-For example, Skywork-Critic-Llama-3.1-70B excels on RewardBench (93.3), yet underperforms on RMB (65.6).
-GPT-4o achieves balanced results, with the second-best RMB score (73.8) and an average of 77.7, but still lags behind leading scalars and reasoning RMs.
-ReasonRMs show clear gains, especially at scale.
-The RM-R1 model family and related ReasonRMs outperform both ScalarRMs and GenRMs once scaled beyond 14B parameters.
-RM-R1-DeepSeek-Distilled-Qwen-32B achieves 83.9 on RM-Bench, the strongest baseline prior to our method, with a solid average of 81.5.
-Similarly, RM-R1-Qwen-Instruct-32B attains 81.2 average.
-These results confirm that structured reasoning pipelines – when scaled – are effective at overcoming shallow analysis and improving sensitivity to subtle reasoning flaws.
-However, smaller reasoning RMs (e.g., JudgeLRM-7B at 64.3 average) trail behind even the weaker ScalarRMs, suggesting that naive reasoning without sufficient scale is not sufficient.
-BR-RM models consistently perform well across all three benchmarks.
-While ScalarRMs dominate RewardBench and ReasonRMs lead RM-Bench, our BR-RMs achieve the best overall balance.
-BR-RM-Qwen-14B reaches 92.1 on RewardBench (close to scalar SOTA), 85.9 on RM-Bench (highest overall), and 74.7 on RMB (top-2 overall), yielding the best average (84.2).
-Even BR-RM-Qwen-8B, with only 8B parameters, surpasses larger baselines such as GPT-4o and INF-ORM-Llama3.1-70B in average score.
-These results highlight adaptive two-turn reasoning not only closes the gap with scalar models on factual benchmarks, but also sets a new bar for reasoning robustness.
+Table 1 compares BR-RMs against state-of-the-art baselines across three benchmarks.
+ScalarRMs plateau on reasoning tasks.
+While models like INF-ORM-Llama3.1-70B dominate RewardBench (95.1), they lag significantly on reasoning-heavy benchmarks (RM-Bench 70.9, RMB 70.5) , suggesting diminishing returns from parameter scaling alone.
+GenRMs trade off consistency.
+Generative models improve robustness but fail to consistently surpass scalars.
+Skywork-Critic excels on RewardBench yet struggles on RMB (65.6), and GPT-4o, despite a balanced profile, still trails leading scalar and reasoning RMs.
+ReasonRMs require scale.
+Reasoning mechanisms show clear gains only beyond 14B parameters.
+The 32B RM-R1 models achieve strong results (up to 83.9 on RM-Bench) , whereas smaller reasoning models (e.g., 7B) underperform even weaker ScalarRMs, indicating that reasoning requires sufficient model capacity.
+BR-RM achieves the best overall balance.
+Our BR-RM-Qwen-14B attains the highest average score (84.2), setting a new SOTA on RM-Bench (85.9) and ranking top-2 on RMB (74.7).
+Notably, our 8B model outperforms significantly larger baselines (e.g., GPT-4o, 70B scalars), confirming that adaptive two-turn reasoning sets a new bar for both efficiency and robustness.
 4.3 Ablation Study
-Table 2: Ablation Study.
-This table summarizes the performance impact of removing key components from our BR-RM models.
-Values in parentheses show the performance drop from the full models, highlighting the contribution of each component.
-Due to space constraints, RewardBench is abbreviated as RB.
-Prompting strategy BR-RM-Qwen-8B BR-RM-Qwen-14B
-RB RMBench RMB Avg RB RMBench RMB Avg
-BR-RM (Full Model) 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
-Branching Only 90.1 82.6 67.4 80.0 90.9 83.2 71.0 81.7
-Unconditioned Rethink 90.6 83.9 69.3 81.3 91.3 84.4 71.9 82.5
-Single-Turn Model 90.4 84.5 70.0 81.6 91.3 85.2 73.3 83.3
+Figure 3: Ablation Study. The performance difference from the full models highlighting the contribution of each component. RewardBench is abbreviated as RB.
 To validate the contributions of our proposed components, we conduct several ablations based on our BR-RM models.
-We systematically remove key architectural components to isolate their impact on performance across our evaluation benchmarks (Table 2).
-Branching Only In this variant, we remove the second “rethinking” turn.
-This ablation caused the most significant performance degradation.
-The first turn is designed for both branching and error detection, which directs its focus toward identifying specific, granular issues.
-Without the second turn, the model’s final judgment relies entirely on this initial pass.
-This can mislead the model to over-emphasize details and minor flaws without having a comprehensive view of the whole response.
-The absence of the holistic re-evaluation step is particularly damaging on nuanced benchmarks like RMB, where the 14B model’s performance drops by a substantial 3.7 points.
-This shows that the second turn is critical for contextualizing initial findings.
-Unconditioned Rethinking Here, we retain the two-turn training but remove branching.
-The first turn goes through all the dimensions while the second turn is prompted with a generic “re-evaluate the response” instruction.
-While this model performs better than the “Branching Only” version, it remained well below our full model.
-For instance, the 14B model sees performance drops of 0.8, 1.5, and 2.8 points on the respective benchmarks.
-This result underscores the importance of adaptive focus.
-Without an explicit directive from the first turn, the second pass is less efficient and prone to the same judgment diffusion it is designed to prevent.
-Single-Turn Model To test the value of our sequential process, we merge the prompts for both turns into a single instruction.
-This approach forces the model to rely on only one-time thinking.
-The resulting prompt is highly complex, demanding that the model handles too many subtasks at the same time: select dimensions, sketch issues, and then re-evaluate to produce a final score.
-This cognitive overload prevents the model from deeply engaging with any single subtask, leading to a shallower analysis.
-The consistent, albeit smaller, performance drop confirms that a deliberate, sequential process is critical for high-quality, reasoned evaluation.
+We systematically remove key architectural components to isolate their impact on performance across our evaluation benchmarks (Figure 3).
+Branching Only Removing the second "rethinking" turn caused the most significant degradation, with a 3.7-point drop on RMB for the 14B model.
+Without the second turn, the model over-emphasizes granular flaws detected in the first pass without a comprehensive perspective.
+This confirms that the holistic re-evaluation step is critical for contextualizing initial findings.
+Unconditioned Rethinking We tested a version with two turns but no branching, using a generic re-evaluation prompt for the second pass.
+Performance dropped by up to 2.8 points, demonstrating that without an explicit directive from the first turn, the second pass suffers from judgment diffusion and reduced efficiency.
+Adaptive focus is essential for localized error correction.
+Single-Turn Model Merging both turns into a single instruction forces the model to perform dimension selection, error sketching, and final scoring simultaneously.
+This cognitive overload leads to shallower analysis compared to our sequential process.
+The resulting performance decline confirms that a deliberate, multi-step pipeline is necessary for high-quality evaluation.
 4.4 Reward Design Analysis
-Table 3: Reward Analysis.
+Figure 4: Reward Analysis.
 Quantitative comparison of our proposed reward design against several ablations.
 Our approach using a direct binary reward consistently achieves the highest scores.
-Due to space constraints, RewardBench is abbreviated as RB.
-Reward BR-RM-Qwen-8B BR-RM-Qwen-14B
-RB RMBench RMB Avg RB RMBench RMB Avg
-Our Reward 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
-Removing Format Checks 90.5 84.2 71.5 82.1 91.3 85.4 73.1 83.3
-Scoring on a Scale 90.0 83.5 70.1 81.2 90.8 85.1 71.7 82.5
-Additional Reward for Branch 86.5 79.0 66.2 77.2 87.8 82.7 69.3 79.9
-To identify the best training strategy, we evaluated several reward designs in Table 3.
-Removing Format Checks First, we tested removing the strict formatting that guides the model’s two-step reasoning process.
-While this simplified the setup, it allowed the model to find shortcuts, such as producing minimal text just to get the final answer right while skipping the intended reasoning steps.
-We found that the format check provides crucial guidance and control.
-It helps the model distinguish between a wrong answer and a badly formatted one; without it, the model would receive the same penalty for incorrect reasoning or just for messy output, which slows down learning.
-Format checks provide essential structure and unambiguous feedback.
-Scoring on a Scale In another experiment, we had the model output a detailed score on a scale, like from -3 to 3.
-This is inspired by the reward dataset, HelpSteer3, whose human-annotated preferences use a similar scale to express preference strength.
-During training, the reward was based on the distance between the predicted score and the ground-truth score.
-To process other datasets with only binary labels, we arbitrarily mapped the winning response to +2 and the losing one to -2.
-This approach failed, possibly due to a disconnect between the training task and the final evaluation: while training rewarded the model for predicting an exact score, our ultimate evaluation only cared about whether the model correctly identified the winning response (a simple binary choice).
-This means that the model had no strong incentive to learn the true meaning behind the scores; it only needed to learn if a score was positive or negative.
-Predicting precise scores is less beneficial when only the winner matters.
-Additional Reward for Branching Finally, we tried to reward the first step of the model’s reasoning process individually using a different reward.
-The key issue here is the absence of a natural reward signal, unlike for the final answer (where ground truth is available).
-Instead, we calculated it from intermediate values generated during the process.
-Specifically, we measured the difference of the rewards with and without branching to quantify the effect of Turn 1.
-But we found that this indirect feedback led to two major practical problems: 1) the approach was computationally expensive, nearly doubling the processing required, and 2) the resulting reward signal was infrequent and erratic, which destabilized training.
-Calculated intermediate rewards created a noisy, unstable signal.
-The Best Reward: Binary Reward with Format Checking.
-After testing these alternatives, we confirmed that the best reward design is the most straightforward one.
-In this setup, the model produces a binary choice (A or B) within a required format and receives a simple reward based on whether it is correct.
-Each of the other designs introduced a flaw – confusing signals, misaligned objectives, or indirect feedback – that hurt performance.
-This shows that a direct alignment between the task and a reward based on the correct answer provides the most effective foundation for training.
-4.5 Training Data Analysis
-Table 4: Training Data Analysis. Impact of training data sources for our BR-RM models. Due to space constraints, RewardBench is abbreviated as RB.
-Model BR-RM-Qwen-8B BR-RM-Qwen-14B
-RB RMBench RMB Avg RB RMBench RMB Avg
-Full training data 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
-w/o HelpSteer3 90.4 83.5 68.0 80.6 91.5 84.4 70.3 82.1
-w/o Skywork 90.6 84.2 66.1 80.3 91.3 85.1 69.7 82.0
-w/o Math-Step-DPO 91.0 84.7 70.7 82.1 91.5 85.0 72.6 83.0
-w/o Code-Preference 91.1 84.8 71.2 82.4 91.8 85.4 74.0 83.7
-To understand each dataset’s contribution, we conducted an ablation study, systematically removing one at a time to measure its impact on model performance.
-From Table 4, this analysis revealed a clear distinction between the roles of foundational and specialized datasets.
-For HelpSteer3, removing this foundational dataset led to a broad performance decline, with a 4.4-point drop on the RMB benchmark and a 1.5-point drop on RMBench.
-This highlights its critical role in establishing general understanding and factual accuracy.
-For Skywork, excluding our only dataset with safety data caused the most severe impact—a 5.0-point drop on the safety-focused RMB benchmark.
-This confirms that explicit safety training is essential for model harmlessness.
-For Math-Step-DPO, removing this specialized dataset predictably weakened reasoning, causing a 2.1-point drop on RMB and a 0.9-point drop on RMBench.
-This shows in-domain data is necessary for complex, multi-step reasoning.
-For Code-Preference, the exclusion of this smaller dataset resulted in the most modest decrease (0.7 points on RMB and 0.5 points on RMBench).
-This suggests its logical reasoning contributions are partially covered by the broader datasets.
+Here RewardBench is abbreviated as RB.
+To identify the best training strategy, we evaluated several reward designs in Figure 4.
+Removing Format Checks Removing strict formatting allowed the model to find shortcuts, skipping reasoning to produce only the final answer.
+Format checks provide essential structure and help the model distinguish between incorrect reasoning and poor formatting.
+Without this control, ambiguous feedback slows down learning.
+Format checks ensure structural integrity and clear, unambiguous feedback.
+Scoring on a Scale We tested a scoring scale to reflect preference strength, mapping binary labels to the range from -3 to 3.
+This failed because the model focused on predicting exact values rather than the relative quality needed to identify a winner.
+This misalignment between the training task and the final evaluation objective hindered performance.
+Predicting precise scores is less beneficial when only the win/loss outcome matters.
+Additional Reward for Branching We attempted to reward the first reasoning step using indirect values calculated from the branching process.
+However, this lacked a natural ground-truth signal and created two problems: computational costs nearly doubled, and the resulting reward signals were too infrequent and erratic to stabilize training.
+Calculated intermediate rewards generate noisy and unstable training signals.
+The Best Reward: Binary Reward with Format Checking The most straightforward design—binary choice within a required format—proved most effective.
+Other designs introduced confusing signals or misaligned objectives that hurt performance.
+Direct alignment between the task and a reward based on the correct answer provides the most effective training foundation.
+4.5 Reasoning Turn Scaling
+Figure 5: Ablation on Turn Scaling (Qwen-14B).
+The trend lines across three benchmarks (RewardBench, RM-Bench, RMB) illustrate that performance consistently peaks at the 2-turn mark and declines with further rethinking, highlighting the diminishing returns of excessive reasoning turns.
+We explore the optimality of the two-turn design by scaling the number of reasoning turns.
+We trained variants of our 14b model with 1, 3, and 4 turns, where additional turns involve iterative rethinking.
+As shown in Figure 5, performance peaks at 2 turns and degrades with additional iterations.
+We hypothesize that beyond two turns, the model begins to “over-think,” hallucinating subtle errors during repeated verification or drifting from the original context.
+This confirms that a single hypothesis generation step followed by a verification step is the most robust configuration, striking an optimal balance between reasoning depth and stability.
+4.6 Distribution of Selected Branches
+Figure 6: Distribution of Selected Branches across RM-Bench Subsets.
+The heatmap visualizes the sparse and domain-specific activation patterns, demonstrating that the model adaptively prioritizes relevant dimensions (e.g., Logical Reasoning for Math, Safety Awareness for Safety) rather than applying a uniform rubric.
+Table 2: Inference Efficiency. Latency is measured per instance on $2\times$ H100 GPUs. Our two-turn 8B model outperforms the larger 32B single-turn baseline in both speed and accuracy.
+Model Type Latency (s) Avg Acc (%)
+InternLM2-7B-Reward Scalar 0.015 72.3
+INF-ORM-Llama3.1-70B Scalar 0.25 78.8
+RM-R1-Qwen-Instruct-7B Gen (1-Turn) 5.1 73.9
+RM-R1-Qwen-Instruct-32B Gen (1-Turn) 10.3 81.2
+BR-RM-Qwen-8B (Ours) Gen (2-Turn) 9.5 82.6
+BR-RM-Qwen-14B (Ours) Gen (2-Turn) 15.9 84.2
+To investigate whether predefined dimensions limit generality, we analyzed branch selection behavior on RM-Bench, observing sparse and adaptive activation (averaging 2.5 dimensions per instance) where specific dimensions like Safety Awareness” or Implementation Capability” heavily dominate their respective domains (91% and 94% activation, respectively).
+This confirms that our universal set functions as a flexible semantic basis rather than a rigid checklist, allowing the model to dynamically map novel failure modes into high-level cognitive categories and effectively prune the search space to prevent focus dilution.
+4.7 Inference Efficiency
+A primary concern with reasoning-based reward models is the increased inference cost compared to traditional scalar models.
+To quantify this trade-off, we conducted a controlled latency benchmark on $2\times$ H100 GPUs.
+While scalar RMs are inherently faster (e.g., InternLM2-7B-Reward takes 0.015s), they hit a distinct performance ceiling, peaking at 78.8% average accuracy with INF-ORM-Llama3.1-70B.
+In the regime of reasoning-based evaluation, our Branch-and-Rethink strategy proves significantly more efficient than simply scaling model parameters.
+As shown in Table 2, ourBR-RM-Qwen-8B requires 9.5s per instance, making it faster than the single-turn baseline RM-R1-Qwen-Instruct-32B (10.3s) while achieving higher average accuracy (82.6% vs.
+81.2%).
+It confirms that allocating test-time compute to a targeted second pass yields better than generating a single broad critique with a larger model.
+Beyond raw latency, the practical feasibility of BR-RM should be viewed through the lens of high-stakes alignment pipelines.
+In offline reinforcement learning settings like Rejection Sampling or PPO, the quality of the reward signal is often the primary bottleneck rather than generation speed; a noisy reward signal from a faster but less accurate scalar model can destabilize training.
+Furthermore, the inference latency of BR-RM is effectively amortized in modern asynchronous RL setups, where reward generation is decoupled from policy updates.
+This architecture prevents the reward model from becoming a blocking factor in the training loop, justifying the investment in “thinking twice” for superior preference labeling.
 5 Conclusion
@@
 Extensions of this work may include integrating verification tools like retrieval and code runners to check claims, automatically generating dynamic evaluation rules to replace static rubrics, and using model uncertainty to adaptively decide when and how deeply to re-evaluate an output.
+Appendix Table of Contents
 Appendix A Preliminary Study on Judgment Diffusion
@@
 The next and most critical step is Token Allocation.
-Here, we prompt a large language model, GPT-4o-0806, to read the generated text from each model and attribute every single sentence to one of the predefined evaluation dimensions (e.g., Information Accuracy, Logical Reasoning).
+Here, we prompt a large language model, GPT-4o-0806 (25), to read the generated text from each model and attribute every single sentence to one of the predefined evaluation dimensions (e.g., Information Accuracy, Logical Reasoning).
 For our method, we combine the tokens from both the initial Adaptive Branching turn and the final Branch-Conditioned Rethinking turn.
@@
   • Scalar RMs: Conventional reward models that directly output a single scalar score without generating intermediate reasoning.
-We compare to SteerLM-RM-70B, Eurus-RM-7B (47), InternLM2-20B-Reward, InternLM2-7B-Reward (6), Skywork-Reward-Gemma-2-27B , Skywork-Reward-Llama-3.1-8B (19), ArmoRM-Llama3-8B-v0.1, Nemotron-4-340B-Reward (1), and INFORM-Llama3.1-70B (9).
+We compare to SteerLM-RM-70B, Eurus-RM-7B (47), InternLM2-20B-Reward, InternLM2-7B-Reward (6), Skywork-Reward-Gemma-2-27B , Skywork-Reward-Llama-3.1-8B (19), ArmoRM-Llama3-8B-v0.1, Nemotron-4-340B-Reward (1), and INF-ORM-Llama3.1-70B (9).
   • Generative RMs (GenRMs): Reward models that produce natural language critiques or rationales before collapsing them into a scalar score.
@@
 Appendix E Detailed Experiment Results
+E.1 Main Results
 Table 4: Results on RewardBench.
@@
 BR-RM-Qwen-14B 97.0 82.4 90.0 98.8 92.1
-In this section, we provide the full experiment results and a more comprehensive coverage of existing baselines. The results of RewardBench, RM-Bench, and RMB are provided in Tables 6, 7, and 8.
 Table 5: The full results of tested reward models on RM-Bench.
@@
 BR-RM-Qwen-14B 0.670 0.810 0.693 0.816 0.747
+In this section, we provide the full experiment results and a more comprehensive coverage of existing baselines. The results of RewardBench, RM-Bench, and RMB are provided in Tables 4, 5, and 6.
+Table 1 presents a comprehensive performance comparison of our proposed models (BR-RMs) against the best-performing baselines across three standard benchmarks.
+ScalarRMs remain competitive but plateau.
+ScalarRMs like INF-ORM-Llama3.1-70B achieve the strongest RewardBench score (95.1) and lead most scalar baselines in average accuracy (78.8).
+However, despite high scores on RewardBench, they lag behind reasoning-focused models on RM-Bench (70.9) and RMB (70.5).
+Smaller ScalarRMs perform relatively well, suggesting simply scaling up parameters without reasoning mechanisms brings diminishing returns.
+GenRMs improve robustness but trade off consistency.
+Generative reward models, which critique before scoring, close some gaps but do not consistently surpass scalar baselines.
+For example, Skywork-Critic-Llama-3.1-70B excels on RewardBench (93.3), yet underperforms on RMB (65.6).
+GPT-4o achieves balanced results, with the second-best RMB score (73.8) and an average of 77.7, but still lags behind leading scalars and reasoning RMs.
+ReasonRMs show clear gains, especially at scale.
+The RM-R1 model family and related ReasonRMs outperform both ScalarRMs and GenRMs once scaled beyond 14B parameters.
+RM-R1-DeepSeek-Distilled-Qwen-32B achieves 83.9 on RM-Bench, the strongest baseline prior to our method, with a solid average of 81.5.
+Similarly, RM-R1-Qwen-Instruct-32B attains 81.2 average.
+These results confirm that structured reasoning pipelines – when scaled – are effective at overcoming shallow analysis and improving sensitivity to subtle reasoning flaws.
+However, smaller reasoning RMs (e.g., JudgeLRM-7B at 64.3 average) trail behind even the weaker ScalarRMs, suggesting that naive reasoning without sufficient scale is not sufficient.
+BR-RM models consistently perform well across all three benchmarks.
+While ScalarRMs dominate RewardBench and ReasonRMs lead RM-Bench, our BR-RMs achieve the best overall balance.
+BR-RM-Qwen-14B reaches 92.1 on RewardBench (close to scalar SOTA), 85.9 on RM-Bench (highest overall), and 74.7 on RMB (top-2 overall), yielding the best average (84.2).
+Even BR-RM-Qwen-8B, with only 8B parameters, surpasses larger baselines such as GPT-4o and INF-ORM-Llama3.1-70B in average score.
+These results highlight adaptive two-turn reasoning not only closes the gap with scalar models on factual benchmarks, but also sets a new bar for reasoning robustness.
+E.2 Ablation Study
+Table 7: Ablation Study.
+This table summarizes the performance impact of removing key components from our BR-RM models.
+Values in parentheses show the performance drop from the full models, highlighting the contribution of each component.
+Due to space constraints, RewardBench is abbreviated as RB.
+Prompting strategy BR-RM-Qwen-8B BR-RM-Qwen-14B
+RB RMBench RMB Avg RB RMBench RMB Avg
+BR-RM (Full Model) 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
+Branching Only 90.1 82.6 67.4 80.0 90.9 83.2 71.0 81.7
+Unconditioned Rethink 90.6 83.9 69.3 81.3 91.3 84.4 71.9 82.5
+Single-Turn Model 90.4 84.5 70.0 81.6 91.3 85.2 73.3 83.3
+E.3 Reward Design Analysis
+To identify the best training strategy, we evaluated several reward designs in Table 8.
+Removing Format Checks First, we tested removing the strict formatting that guides the model’s two-step reasoning process.
+While this simplified the setup, it allowed the model to find shortcuts, such as producing minimal text just to get the final answer right while skipping the intended reasoning steps.
+We found that the format check provides crucial guidance and control.
+It helps the model distinguish between a wrong answer and a badly formatted one; without it, the model would receive the same penalty for incorrect reasoning or just for messy output, which slows down learning.
+Format checks provide essential structure and unambiguous feedback.
+Scoring on a Scale In another experiment, we had the model output a detailed score on a scale, like from -3 to 3.
+This is inspired by the reward dataset, HelpSteer3, whose human-annotated preferences use a similar scale to express preference strength.
+During training, the reward was based on the distance between the predicted score and the ground-truth score.
+To process other datasets with only binary labels, we arbitrarily mapped the winning response to +2 and the losing one to -2.
+This approach failed, possibly due to a disconnect between the training task and the final evaluation: while training rewarded the model for predicting an exact score, our ultimate evaluation only cared about whether the model correctly identified the winning response.
+This means that the model had no strong incentive to learn the true meaning behind the scores; it only needed to learn if a score was positive or negative.
+Predicting precise scores is less beneficial when only the winner matters.
+Additional Reward for Branching Finally, we tried to reward the first step of the model’s reasoning process individually using a different reward.
+The key issue here is the absence of a natural reward signal, unlike for the final answer (where ground truth is available).
+Instead, we calculated it from intermediate values generated during the process.
+Specifically, we measured the difference of the rewards with and without branching to quantify the effect of Turn 1.
+But we found that this indirect feedback led to two major practical problems: 1) the approach was computationally expensive, nearly doubling the processing required, and 2) the resulting reward signal was infrequent and erratic, which destabilized training.
+Calculated intermediate rewards created a noisy, unstable signal.
+The Best Reward: Binary Reward with Format Checking.
+After testing these alternatives, we confirmed that the best reward design is the most straightforward.
+In this setup, the model produces a binary choice within a required format and receives a simple reward based on whether it is correct.
+Each of the other designs introduced a flaw – confusing signals, misaligned objectives, or indirect feedback – that hurt performance.
+This shows that a direct alignment between the task and a reward based on the correct answer provides the most effective foundation for training.
+Table 8: Reward Analysis.
+Quantitative comparison of our proposed reward design against several ablations.
+Our approach using a direct binary reward consistently achieves the highest scores.
+Due to space constraints, RewardBench is abbreviated as RB.
+Reward BR-RM-Qwen-8B BR-RM-Qwen-14B
+RB RMBench RMB Avg RB RMBench RMB Avg
+Our Reward 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
+Removing Format Checks 90.5 84.2 71.5 82.1 91.3 85.4 73.1 83.3
+Scoring on a Scale 90.0 83.5 70.1 81.2 90.8 85.1 71.7 82.5
+Additional Reward for Branch 86.5 79.0 66.2 77.2 87.8 82.7 69.3 79.9
+E.4 Training Data Analysis
+To understand each dataset’s contribution, we conducted an ablation study, systematically removing one at a time to measure its impact on model performance.
+From Table 9, this analysis revealed a clear distinction between the roles of foundational and specialized datasets.
+For HelpSteer3, removing this foundational dataset led to a broad performance decline, with a 4.4-point drop on the RMB benchmark and a 1.5-point drop on RMBench.
+This highlights its critical role in establishing general understanding and factual accuracy.
+For Skywork, excluding our only dataset with safety data caused the most severe impact—a 5.0-point drop on the safety-focused RMB benchmark.
+This confirms that explicit safety training is essential for model harmlessness.
+For Math-Step-DPO, removing this specialized dataset predictably weakened reasoning, causing a 2.1-point drop on RMB and a 0.9-point drop on RMBench.
+This shows in-domain data is necessary for complex reasoning.
+For Code-Preference, the exclusion of this smaller dataset resulted in the most modest decrease (0.7 points on RMB and 0.5 points on RMBench).
+This suggests its logical reasoning contributions are partially covered by the broader datasets.
+Table 9: Training Data Analysis. Impact of training data sources for our BR-RM models. Due to space constraints, RewardBench is abbreviated as RB.
+Model BR-RM-Qwen-8B BR-RM-Qwen-14B
+RB RMBench RMB Avg RB RMBench RMB Avg
+Full training data 91.0 85.0 71.8 82.6 92.1 85.9 74.7 84.2
+  w/o HelpSteer3 90.4 83.5 68.0 80.6 91.5 84.4 70.3 82.1
+  w/o Skywork 90.6 84.2 66.1 80.3 91.3 85.1 69.7 82.0
+  w/o Math-Step-DPO 91.0 84.7 70.7 82.1 91.5 85.0 72.6 83.0
+  w/o Code-Preference 91.1 84.8 71.2 82.4 91.8 85.4 74.0 83.7
 Appendix F Case Study